# 3cm3cm3cm3cm

### v1

@@ 3.1. Column determinants and Berezinians
Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$. The parity of a homogeneous element $a\in\mathcal{A}$ is denoted by $|a|$, which lies in ${\mathbb{Z}}_{2}$. Fix $k\in{\mathbb{N}}$. For any $k\times k$ matrix $A=\big{[}a_{i,j}\big{]}_{i,j=1,\ldots,k}$ over $\mathcal{A}$, the column determinant of $A$ is defined to be
${\rm cdet}\,A=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{\sigma(1),1}\ldots a_{\sigma(k),k}.$
Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$ and $l(\sigma)$ denotes the length of $\sigma$. For any nonempty subset $P=\{i_{1}<\ldots<i_{\ell}\}$ of $\{1,\ldots,k\}$, the matrix $A_{P}:=\big{[}a_{i,j}\big{]}_{i,j\in P}$ is called a standard submatrix of $A$.
Assume that $A$ has a two-sided inverse $A^{-1}=\big{[}\widetilde{a}_{i,j}\big{]}$. For all $i,j=1,\ldots,k$, the $(i,j)$th quasideterminant of $A$ is defined to be $|A|_{ij}:=\widetilde{a}_{j,i}^{-1}$ provided that $\widetilde{a}_{j,i}$ has an inverse in $\mathcal{A}$. According to [GGRW, Proposition 1.2.6], if $k\geq 2$, then
$|A|_{ij}=a_{i,j}-r_{i}^{j}(A^{ij})^{-1}c_{j}^{i},$
where $A^{ij}$ is the $(k-1)\times(k-1)$ submatrix of $A$ obtained by deleting the $i$th row and the $j$th column of $A$ and is assumed to be invertible, $r_{i}^{j}$ is the $i$th row of $A$ with $a_{i,j}$ removed, and $c_{j}^{i}$ is the $j$th column of $A$ with $a_{i,j}$ removed. Following the notation of [GGRW], it is convenient to write
@@
$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix},$
@@
A $k\times k$ matrix $A$ over $\mathcal{A}$ is called sufficiently invertible if every principal quasiminor of $A$ is well defined, and $A$ is called amply invertible if each of its standard submatrices is sufficiently invertible.
Let ${{\bf s}}=(s_{1},\ldots,s_{m+n})$ be a sequence of 0’s and 1’s such that exactly $m$ of the $s_{i}$’s are 0 and the others are 1. We call such a sequence a $0^{m}1^{n}$-sequence. Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 00’s, followed by $n_{1}$ copies of $1$’s, and so on. The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
Let ${\bf s}\in\mathcal{S}_{{m|n}}$. For any $\sigma\in\mathfrak{S}_{m+n}$ and any $(m+n)\times(m+n)$ matrix $A=[a_{i,j}]$ over $\mathcal{A}$, we define ${\bf s}^{\sigma}=\left(s_{\sigma^{-1}(1)},s_{\sigma^{-1}(2)},\ldots,s_{\sigma^{-1}(m+n)}\right)$ and $A^{\sigma}=\big{[}a_{\sigma^{-1}(i),\sigma^{-1}(j)}\big{]}$. We say that $A$ is of type ${\bf s}$ if $a_{i,j}$ is a homogeneous element of parity $|a_{i,j}|=\bar{s}_{i}+\bar{s}_{j}$ for any $i,j=1,\ldots,m+n$.
For any $(m+n)\times(m+n)$ sufficiently invertible matrix $A$ of type ${\bf s}$ over $\mathcal{A}$, the Berezinian of type ${\bf s}$ of $A$ is defined to be
${\rm Ber}^{\bf s}\,A=d_{1}(A)^{\hat{s}_{1}}\ldots d_{m+n}(A)^{\hat{s}_{m+n}},$
where $\hat{s}_{i}:=(-1)^{s_{i}}$ for $1\leq i\leq m+n$ (see [HM, (3.3)]). We refer the reader to [Ber, Na, MR] for earlier definitions of Berezinians. We have the following proposition.
@@
Let $A$ be an $(m+n)\times(m+n)$ amply invertible matrix of type ${\bf s}$ over $\mathcal{A}$. Fix $k\in\{1,\ldots,m+n-1\}$. We write
@@
${\rm Ber}^{\bf s}\,A={\rm Ber}^{{\bf s}^{\prime}}W\cdot{\rm Ber}^{{\bf s}^{\prime\prime}}\big{(}Z-YW^{-1}X\big{)}.$
An $(m+n)\times(m+n)$ matrix $A=\big{[}a_{i,j}\big{]}$ over $\mathcal{A}$ is called a Manin matrix of type ${\bf s}$ if $A$ is a matrix of type ${\bf s}$ satisfying the following relations
$[a_{i,j},a_{k,l}]=(-1)^{s_{i}s_{j}+s_{i}s_{k}+s_{j}s_{k}}[a_{k,j},a_{i,l}]\qquad\text{for all}\quad i,j,k,l=1,\ldots,m+n,$
where $[a,b]:=ab-(-1)^{|a||b|}ba$ for any homogeneous elements $a,b\in\mathcal{A}$. We recall some basic facts about Manin matrices.
@@
Let $A$ be an $(m+n)\times(m+n)$ Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Then
@@
    $A^{-1}$ is an $(m+n)\times(m+n)$ Manin matrix of type ${\bf s}$ over $\mathcal{A}$.
@@
    If $P=\{i_{1}<\ldots<i_{\ell}\}$ is a nonempty subset of $\{1,\ldots,m+n\}$, then the standard submatrix $A_{P}$ of $A$ is a Manin matrix of type ${{\bf s}}_{P}:=(s_{i_{1}},\ldots s_{i_{\ell}})$.
  3. (iii)
@@
The following proposition asserts that the Berezinian of a Manin matrix is invariant under the action of a permutation.
@@
Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Then
$\hskip 28.45274pt{\rm Ber}^{{\bf s}^{\sigma}}A^{\sigma}={\rm Ber}^{\bf s}\,A\qquad\mbox{for any $\sigma\in\mathfrak{S}_{m+n}$.}$
We may decompose ${\rm Ber}^{\bf s}\,A$ in Proposition 3.1 in a different way if $A$ is a Manin matrix.
@@
Let $A$ be an $(m+n)\times(m+n)$ amply invertible Manin matrix of type ${\bf s}$ over $\mathcal{A}$. Write $A$ as in (3.1). Then $W-XZ^{-1}Y$ and $Z$ are sufficiently invertible Manin matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}\,A={\rm Ber}^{{\bf s}^{\prime\prime}}Z\cdot{\rm Ber}^{{\bf s}^{\prime}}\big{(}W-XZ^{-1}Y\big{)}.$ (3.2)
@@
The first assertion is clear from Proposition 3.2. Let $\sigma\in\mathfrak{S}_{m+n}$ be the permutation defined by $\sigma(k+i)=i$ for $i=1,\ldots,m+n-k$ and $\sigma(j)=m+n-k+j$ for $j=1,\ldots,k$. Then
@@
Proposition 3.4 will be used in Section 4.2. We also have the following proposition.
@@
Let ${\bf s}=(0^{m})$. For any $m\times m$ sufficiently invertible Manin matrix $A$ of type ${\bf s}$ over $\mathcal{A}$, we have ${\rm Ber}^{\bf s}\,A={\rm cdet}\,\,A$.
3.2. Pseudo-differential operators
Let $\mathcal{A}$ be an associative unital superalgebra over ${\mathbb{C}}$ and $z$ an even variable. We denote by $\mathcal{A}[\kern-1.1pt[z]\kern-1.1pt]$ and $\mathcal{A}(\!(z)\!)$ the superalgebras of formal power series and Laurent series in $z$ with coefficients in $\mathcal{A}$, respectively. Let $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ denote the set of all formal series of the form
$\sum_{\ell=-\infty}^{s}\sum_{k=-\infty}^{r}a_{k\ell}z^{k}\partial_{z}^{\ell},$
where $r,s\in{\mathbb{Z}}$ and $a_{k\ell}\in\mathcal{A}$. We may endow $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ with a superalgebra structure using the rules:
$\partial_{z}\partial_{z}^{-1}=\partial_{z}^{-1}\partial_{z}=1,\quad\partial_{z}^{k}z^{\ell}=\sum_{i=0}^{\infty}\binom{k}{i}\binom{\ell}{i}i!\,z^{\ell-i}\,\partial_{z}^{k-i},\quad\mbox{for $k,\ell\in{\mathbb{Z}}$.}$ (3.3)
Here, e.g., ${\binom{k}{i}:=\frac{k(k-1)\ldots(k-i+1)}{i!}}$. The superalgebra $\mathcal{A}(\!(z^{-1},\partial_{z}^{-1})\!)$ is called the superalgebra of pseudo-differential operators over $\mathcal{A}$.
$\omega(az^{k}\partial_{z}^{\ell})=\omega(a)z^{\ell}\partial_{z}^{k}$ (3.4)
for $a\in\mathcal{A}$ and $k,\ell\in{\mathbb{Z}}$.
Proposition 3.6.
Proof.
Clearly, $\omega^{2}=1$. It remains to show that $\omega$ is an anti-homomorphism. For $a,b\in\mathcal{A}$ and $k,\ell,r,s\in{\mathbb{Z}}$, we have
$\omega((az^{k}\partial_{z}^{\ell})(bz^{r}\partial_{z}^{s}))$ $=$ $\omega(abz^{k}\partial_{z}^{\ell}z^{r}\partial_{z}^{s})$
$=$ $\sum_{i=0}^{\infty}\binom{\ell}{i}\binom{r}{i}i!\,\omega(abz^{k+r-i}\partial_{z}^{\ell+s-i})$
$=$ $\sum_{i=0}^{\infty}\binom{\ell}{i}\binom{r}{i}i!\,\omega(ab)z^{\ell+s-i}\partial_{z}^{k+r-i}$
$=$ $\omega(b)\omega(a)z^{s}\partial_{z}^{r}z^{\ell}\partial_{z}^{k}$
$=$ $\left(\omega(b)z^{s}\partial_{z}^{r}\right)\left(\omega(a)z^{\ell}\partial_{z}^{k}\right)$
$=$ $\omega(bz^{r}\partial_{z}^{s})\,\omega(az^{k}\partial_{z}^{\ell}),$
and the proposition follows. ∎
3.3. The Bethe algebra for ${\mathfrak{gl}}_{d}$ and Frobenius algebras
Let
${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{\ell}.$
Define $\mu_{\bf w}\in{\mathfrak{gl}}_{d}^{*}$ by
$\hskip 28.45274pt\mu_{\bf w}(e_{ab})=-\delta_{a,b}w_{a}\qquad\mbox{for $1\leq a\leq d$.}$
The choice of the minus sign will be clear below.
Recall the map $\Psi^{\mu_{\bf w}}$ defined in (2.3). We write
$\Psi^{\bf w}=\Psi^{\mu_{\bf w}}.$
@@ 3.3. The Bethe algebra for ${\mathfrak{gl}}_{d}$ and Frobenius algebras
Here $t$ and $\partial_{t}$ satisfy the rules similar to (3.3). The map $\Psi^{\bf w}$ extends to an algebra homomorphism
@@
such that $\Psi^{\bf w}(\tau)=\partial_{z}$. Let
@@
which is a $d\times d$ Manin matrix. We have
${\rm cdet}\,(\mathcal{T}_{d})=\tau^{d}+\sum_{i=1}^{d}S_{i}\tau^{i-1}.$ (3.5)
By [CM, Theorem 3.1] (see also [CT]), $S_{1},\ldots,S_{d}$ form a complete set of Segal–Sugawara vectors for the Feigin–Frenkel center ${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$. That is, the set $\{{\rm T}^{r}S_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}\}$ is algebraically independent, and
${\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})={\mathbb{C}}[{\rm T}^{r}S_{i}\,|\,i=1,\ldots,d,\,r\in{\mathbb{Z}}_{+}].$
Now consider the $d\times d$ Manin matrix
${\mathcal{L}}_{d}^{\bf w}:=\Big{[}\Psi^{\bf w}\big{(}\delta_{a,b}\tau+e_{ab}[-1]\big{)}\Big{]}_{a,b=1,\ldots,d}=\Big{[}\delta_{a,b}(\partial_{z}-w_{a})-e_{ab}(z)\Big{]}_{a,b=1,\ldots,d}.$
@@
where $S_{i}^{{\bf w}}(z):=\Psi^{\bf w}(S_{i})$.
The following is well-known and matches the definition of the Bethe algebra in [MTV1].
Proposition 3.7.
The Bethe algebra $\mathcal{B}_{d}^{\bf w}$ is generated by the coefficients of $S_{i}^{{\bf w}}(z)$ for $i=1,\ldots,d$.
A Frobenius algebra $\mathcal{A}$ is a finite-dimensional commutative associative unital algebra which admits a nondegenerate symmetric bilinear form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) such that
Recall the notations (2.1) and (2.2). Let $V_{1},\ldots,V_{\ell}$ be finite-dimensional irreducible ${\mathfrak{gl}}_{d}$-modules, and let ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$. As ${\mathfrak{h}}_{d}\subseteq{\mathfrak{gl}}_{d}^{\mu_{\bf w}}$, the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ acts on any weight space of $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$. We will explain why $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra for ${\bf w}\in{\boldsymbol{X}}_{\!d}$.
Define an anti-involution $\boldsymbol{\theta}:{\mathfrak{gl}}_{d}\longrightarrow{\mathfrak{gl}}_{d}$ by
$\boldsymbol{\theta}(e_{ab})=e_{ba}\qquad\mbox{for $1\leq a,b\leq d$.}$ (3.7)
The map $\boldsymbol{\theta}$ is called the Cartan anti-involution on ${\mathfrak{gl}}_{d}$.
For $i=1,\ldots,\ell$, $V_{i}$ is an irreducible highest weight ${\mathfrak{gl}}_{d}$-module with a dominant integral highest weight. Let $v_{i}$ be a highest weight vector of $V_{i}$. There is a unique nondegenerate symmetric bilinear form $(\cdot,\cdot)_{i}$ on $V_{i}$ defined by
$(v_{i},v_{i})_{i}=1\quad\text{and}\quad(Av,v^{\prime})_{i}=(v,\boldsymbol{\theta}(A)v^{\prime})_{i}$
for all $v,v^{\prime}\in V_{i}$ and $A\in{\mathfrak{gl}}_{d}$ ([Sh]). The bilinear form $(\cdot,\cdot)_{i}$ is called the Shapovalov form on $V_{i}$. It is easy to see that if $v$ and $v^{\prime}$ are weight vectors of $V_{i}$ of distinct weights, then $(v,v^{\prime})_{i}=0$.
Let (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denote the tensor form of the bilinear forms $(\cdot,\cdot)_{1},\ldots,(\cdot,\cdot)_{\ell}$, which is called the tensor Shapovalov form. It is a nondegenerate symmetric bilinear form on $\underline{V\mkern-4.0mu}\mkern 4.0mu$. Let ${\bf w}\in{\mathbb{C}}^{d}$. By [MTV1, Theorem 9.1] and [Lu, Lemma 2.6], the Bethe algebra $\mathcal{B}_{d}^{\bf w}$ is symmetric with respect to (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ), i.e.,
$\hskip 42.67912pt({\boldsymbol{b}}v,v^{\prime})=(v,{\boldsymbol{b}}v^{\prime})\quad\mbox{for ${\boldsymbol{b}}\in\mathcal{B}_{d}^{\bf w}$ and $v,v^{\prime}\in\underline{V\mkern-4.0mu}\mkern 4.0mu$}.$
Theorem 3.8.
For any ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the Bethe algebra $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a Frobenius algebra. Moreover, $(\mathcal{B}_{d}^{\bf w})_{{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}}$ is a Frobenius algebra for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
Proof.
The first statement follows from the above discussion on the tensor Shapovalov form (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on $\underline{V\mkern-4.0mu}\mkern 4.0mu$ in view of Theorem 2.1 and [Lu, Lemma 2.7]. As (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) restricts to a nondegenerate symmetric bilinear form on ${\underline{V\mkern-4.0mu}\mkern 4.0mu}_{\!\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$, the second statement follows. ∎
The following is a consequence of [Lu, Lemma 1.3].
Theorem 3.9.
For any ${\bf w}\in{\boldsymbol{X}}_{\!d}$ and ${\bf z}\in{\boldsymbol{X}}_{\!\ell}$, the following properties hold:
  1. (i)
    The Bethe algebra $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a maximal commutative subalgebra of ${\rm End}(\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z}))$ of dimension $\dim\!\left({\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}\right)$.
  2. (ii)
    Every eigenspace of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is one-dimensional, and the set of eigenspaces of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is in bijective correspondence with the set of maximal ideals of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$.
  3. (iii)
    Every generalized eigenspace of $(\mathcal{B}_{d}^{\bf w})_{\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}$ is a cyclic $\mathcal{B}_{d}^{\bf w}$-module.
Remark 3.10.
Each statement of Theorem 3.9 holds if $\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})$ is replaced with ${\underline{V\mkern-4.0mu}\mkern 4.0mu({\bf z})}_{\gamma}$ for any weight $\gamma$ of $\underline{V\mkern-4.0mu}\mkern 4.0mu$.
3.4. The Bethe algebra for ${\mathfrak{gl}}_{p+m|q+n}$
@@
It is straightforward to verify that $\mathcal{T}_{p+m|q+n}$ is an amply invertible $(p+q+m+n)\times(p+q+m+n)$ Manin matrix of type ${\bf s}$ over $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)$. The verification is analogous to [MR, Lemma 3.1]. The following is an immediate consequence of Proposition 3.3.
Proposition 3.11.
For each $\sigma\in\mathfrak{S}_{p+q+m+n}$,
${\rm Ber}\,^{{\bf s}^{\sigma}}\!\left(\mathcal{T}_{p+m|q+n}^{\sigma}\right)={\rm Ber}^{\bf s}\,\!\left(\mathcal{T}_{p+m|q+n}\right).$
We have the expansion

### v2

@@ 4.1. Column determinants and Berezinians
Let us review the notions of column determinants and Berezinians.
Let $\mathcal{A}$ be an associative unital superalgebra. The parity of a homogeneous element $a\in\mathcal{A}$ is denoted by $|a|$, which lies in ${\mathbb{Z}}_{2}$. Fix $k\in{\mathbb{N}}$. For any $k\times k$ matrix $A=\big[a_{i,j}\big]_{i,j=1,\ldots,k}$ over $\mathcal{A}$, the column determinant of $A$ is defined to be
${\rm cdet}(A)=\sum_{\sigma\in\mathfrak{S}_{k}}(-1)^{l(\sigma)}\,a_{\sigma(1),1}\ldots a_{\sigma(k),k}.$
Here $\mathfrak{S}_{k}$ denotes the symmetric group on $\{1,\ldots,k\}$ and $l(\sigma)$ denotes the length of $\sigma$.
Assume that $A$ has a two-sided inverse $A^{-1}=\big[\widetilde{a}_{i,j}\big]$. For all $i,j=1,\ldots,k$, the $(i,j)$th quasideterminant of $A$ is defined to be $|A|_{ij}:=\widetilde{a}_{j,i}^{-1}$ provided that $\widetilde{a}_{j,i}$ has an inverse in $\mathcal{A}$. Following the notation of [GGRW], it is convenient to write
@@
$d_{i}(A)=\begin{vmatrix}a_{1,1}&\ldots&a_{1,i}\\ \ldots&\ldots&\ldots\\ a_{i,1}&\ldots&\framebox{$a_{i,i}$}\end{vmatrix}$,\
@@
Let $A=\big[a_{i,j}\big]_{i,j=1,\ldots,k}$ be a $k\times k$ matrix over $\mathcal{A}$. For any nonempty subset $P=\{i_{1}<\ldots<i_{p}\}$ of $\{1,\ldots,k\}$, the matrix $A_{P}:=\big[a_{i,j}\big]_{i,j\in P}$ is called a standard submatrix of $A$. We say that $A$ is sufficiently invertible if every principal quasiminor of $A$ is well defined, and that $A$ is amply invertible if each of its standard submatrices is sufficiently invertible.
A sequence $(s_{1},\ldots,s_{m+n})$ of 0’s and 1’s is called a $0^{m}1^{n}$-sequence if exactly $m$ of the $s_{i}$’s are 0 and the others are 1. Every $0^{m}1^{n}$-sequence can be written in the form $(0^{m_{1}},1^{n_{1}},\ldots,0^{m_{r}},1^{n_{r}}),$ where the sequence begins with $m_{1}$ copies of 0’s, followed by $n_{1}$ copies of $1$’s, and so on. The set of all $0^{m}1^{n}$-sequences is denoted by $\mathcal{S}_{{m|n}}$.
Fix ${\bf s}:=(s_{1},\ldots,s_{m+n})\in\mathcal{S}_{{m|n}}$. For any $\sigma\in\mathfrak{S}_{m+n}$ and any $(m+n)\times(m+n)$ matrix $A=[a_{i,j}]_{i=1,\ldots,m+n}$ over $\mathcal{A}$, we define ${\bf s}^{\sigma}=\left(s_{\sigma^{-1}(1)},s_{\sigma^{-1}(2)},\ldots,s_{\sigma^{-1}(m+n)}\right)$ and $A^{\sigma}=\big[a_{\sigma^{-1}(i),\sigma^{-1}(j)}\big]_{i=1,\ldots,m+n}$. We say that $A$ is of type $\bf s$} if $a_{i,j}$ is homogeneous of parity $|a_{i,j}|=\bar{s}_{i}+\bar{s}_{j}$ for any $i,j=1,\ldots,m+n$.
For any sufficiently invertible matrix $A$ of type $\bf s$} over $\mathcal{A}$, the Berezinian of type $\bf s$} of $A$ is defined to be
${\rm Ber}^{\bf s}(A)=d_{1}(A)^{\hat{s}_{1}}\ldots d_{m+n}(A)^{\hat{s}_{m+n}},$
where $\hat{s}_{i}:=(-1)^{s_{i}}$ for $i=1,\ldots,m+n$ (see [HM, (3.3)]). We refer the reader to [Ber, Na, MR] for earlier definitions of Berezinians. The following proposition is useful.
@@
Let $A$ be an amply invertible matrix of type $\bf s$} over $\mathcal{A}$. Fix $k\in\{1,\ldots,m+n-1\}$. We write
@@
${\rm Ber}^{\bf s}(A)={\rm Ber}^{{\bf s}^{\prime}}(W)\cdot{\rm Ber}^{{\bf s}^{\prime\prime}}\big(Z-YW^{-1}X\big).$
A matrix $A=[a_{i,j}]_{i=1,\ldots,m+n}$ over $\mathcal{A}$ is called a Manin matrix of type $\bf s$} if $A$ is a matrix of type $\bf s$} satisfying the following relations
$a_{i,j},a_{k,l}]=(-1)^{s_{i}s_{j}+s_{i}s_{k}+s_{j}s_{k}}[a_{k,j},a_{i,l}]\qquad\text{for all}\quad i,j,k,l=1,\ldots,m+n,$
where $a,b]:=ab-(-1)^{|a||b|}ba$ for any homogeneous elements $a,b\in\mathcal{A}$. We recall some basic facts about Manin matrices.
@@
Let $A$ be a Manin matrix of type $\bf s$} over $\mathcal{A}$. Then
@@
    If $P=\{i_{1}<\ldots<i_{p}\}$ is a nonempty subset of $\{1,\ldots,m+n\}$, then the standard submatrix $A_{P}$ is a Manin matrix of type $(s_{i_{1}},\ldots s_{i_{p}})$.
@@
The following proposition asserts that the Berezinian of a Manin matrix is invariant under any permutation.
@@
Let $A$ be an amply invertible Manin matrix of type $\bf s$} over $\mathcal{A}$. Then
$\hskip 28.45274pt{\rm Ber}^{{\bf s}^{\sigma}}(A^{\sigma})={\rm Ber}^{\bf s}(A)\qquad\mbox{for any $\sigma\in\mathfrak{S}_{m+n}$.}$
We may decompose ${\rm Ber}^{\bf s}(A)$ in Proposition 4.1 in a different way if $A$ is a Manin matrix.
@@
Let $A$ be an amply invertible Manin matrix of type $\bf s$} over $\mathcal{A}$. Write $A$ as in (4.1). Then $W-XZ^{-1}Y$ and $Z$ are sufficiently invertible matrices of types ${\bf s}^{\prime}:=(s_{1},\ldots,s_{k})$ and ${\bf s}^{\prime\prime}:=(s_{k+1},\ldots,s_{m+n})$, respectively. Moreover,
${\rm Ber}^{\bf s}(A)={\rm Ber}^{{\bf s}^{\prime\prime}}(Z)\cdot{\rm Ber}^{{\bf s}^{\prime}}\big(W-XZ^{-1}Y\big).$ (4.2)
@@
The first assertion is clear. Let $\sigma\in\mathfrak{S}_{m+n}$ be the permutation defined by $\sigma(k+i)=i$ for $i=1,\ldots,m+n-k$ and $\sigma(j)=m+n-k+j$ for $j=1,\ldots,k$. Then
@@
We also have the following.
@@
Let $A$ be a sufficiently invertible Manin matrix of type ${\bf s}_{0}:=(0^{m})$ over $\mathcal{A}$. Then ${\rm Ber}^{{\bf s}_{0}}(A)={\rm cdet}(A)$.
4.2. The Bethe algebras for ${\mathfrak{gl}}_{d}$ and ${\mathfrak{gl}}_{p+m|q+n}$
@@
Here $t$ and $\partial_{t}$ satisfy the rules similar to (2.8). Let ${\bf w}=(w_{1},\ldots,w_{d})\in{\mathbb{C}}^{d}$, and recall $\mu_{\bf w}\in{\mathfrak{gl}}_{d}^{*}$ defined by (3.12) and (3.13). Write
$\Psi^{\bf w}=\Psi^{\mu_{\bf w}}.$
The map $\Psi^{\bf w}$ extends to an algebra homomorphism
@@
such that $\Psi^{\bf w}(\tau)=\partial_{z}$.
Let
@@
which is a Manin matrix of type $(0^{d})$. The vectors $S_{1},\ldots,S_{d}\in{\mathfrak{z}}(\widehat{{\mathfrak{gl}}}_{d})$, mentioned in (3.2), are the coefficients of ${\rm cdet}(\mathcal{T}_{d})\in U(t^{-1}{\mathfrak{gl}}_{d}[t^{-1}])[\tau]$; more precisely,
${\rm cdet}(\mathcal{T}_{d})=\tau^{d}+\sum_{i=1}^{d}S_{i}\tau^{i-1}$ (4.3)
(see [CM, Theorem 3.1]). Now consider the Manin matrix
${\mathcal{L}}_{d}^{\bf w}:=\Big[\Psi^{\bf w}\big(\delta_{a,b}\tau+e_{ab}[-1]\big)\Big]_{a,b=1,\ldots,d}=\Big[\delta_{a,b}(\partial_{z}-w_{a})-e_{ab}(z)\Big]_{a,b=1,\ldots,d}.$ (4.4)
@@
where $S_{i}^{{\bf w}}(z):=\Psi^{\bf w}(S_{i})$. Write
$\mathcal{B}_{d}^{\bf w}=\mathcal{B}_{{\mathfrak{gl}}_{d}}^{\bf w}.$
Recall from Proposition 3.8 that $\mathcal{B}_{d}^{\bf w}$ is equal to the subalgebra of $U({\mathfrak{gl}}_{d}[t])$ generated by the coefficients of $S_{i}^{{\bf w}}(z)$ for $i=1,\ldots,d$.
@@
It is straightforward to verify that $\mathcal{T}_{p+m|q+n}$ is an amply invertible Manin matrix of type $\bf s$} over $U(t^{-1}{\mathfrak{gl}}_{p+m|q+n}[t^{-1}])(\!(\tau^{-1})\!)$ (cf. [MR, Lemma 3.1]). We have the following expansion: